A 2.1kg mass is on a frictionless 25 degree inclined plane. An 8.0 N horizontal force is applied to the mass. Calculate the time it takes for the mass to move 0.30m on the plane from rest.

a) 0.30s

b) 0.55s

c) 0.93s
Answer #1

answer) we have the sum of the forces in x direction and from newtons law

Fx=ma

Fg-Fh=ma

Fg=mgsin ( gravity force)

Fh= 8cos

so now we have

mgsin-Fcos=m*a

2.1*9.8*sin25-8*cos25=2.1*a

a=0.689 m/s2

now from kinematics

s=ut+1/2at2

u=0 , since we start from rest

s=1/2at2

0.30=1/2*0.689*t2

t=2*0.3/0.689

t=0.93 s

so answer is c) 0.93 s.